# Frontier AI’s Impact on the Cybersecurity Landscape

### v2

@@ 7.2 Enhance Cyber Defenses
Improve proactive testing. First, more efforts are needed in designing automated penetration testing benchmarks and agent-based tools. Regarding the agent, future efforts could focus on providing richer tool support, such as searching for new vulnerabilities in external databases or the internet and tools that help maintain access. With rich tools, we also need to design agents with stronger planning and more capabilities beyond the current ones, such as privilege escalation, compromising domain controllers, and lateral movements.
@@
Second, for vulnerability detection, we recommend leveraging frontier AI to solve the key challenges of existing program analysis-based methods. For static analysis, promising directions include AI-based state pruning to alleviate state explosion and extracting static rules through model explanation to enhance knowledge bases. For fuzzing, one can explore efficient seed/mutator scheduling with RL and LLM-based grammar generation for domain-specific fuzzers. Third, we recommend fine-tuning customized models for better program state understanding and vulnerability reasoning. This requires high-quality training data with correct labels and more program artifacts and context, especially dynamic information, to teach the model program states [81]. We then need to design customized fine-tuning algorithms to gain testing-phase reasoning capabilities. Distillation of large reasoning models through supervised fine-tuning and reinforcement learning with proper reward functions are promising directions, which are effective in code generation and math problem-solving [168, 203, 112]. It is critical to scale AI-driven vulnerability detections beyond function levels to file and repository levels. Fine-tuning models with long context limits and building AI agents with program analysis tools (e.g., AST analysis and data/control flow analysis) could be promising directions.
Finally, we recommend constructing vulnerability detection benchmarks that avoid data leakages in the testing set and have balanced vulnerability distributions that consider the long tails. Having human experts work with AI can potentially strike a balance between reducing manual efforts and improving data quality.
@@
Enhance attack detection. We first recommend constructing high-quality benchmarks that minimize the critical data and label issues, such as duplicated data, spurious correlations, biased samples, and wrong and missing labels. A recent research [18] provided a comprehensive summary of these issues. Given the challenges of collecting high-quality real-world attack data, collecting data from controlled environments [29] and creating synthetic data [302] could be efficient ways. It is necessary to construct hidden test sets that include unseen attacks. These hidden sets will also help continuously retrain AI-driven defenses (Once the currently hidden test sets are used for training, new test sets must be constructed and kept separate from the training data). Such practices have been applied in the natural language field: hiding test sets (e.g. SQuAD [228]), updating test sets with recent data (e.g. LiveBench [294]), or estimating and resolving training-test overlap [328].
@@
Regarding the learning models and methods, although transformers address key challenges of other DNN models, it is still important to develop customized learning methods for security-specific challenges, e.g., attack drifting and extremely imbalanced classes, as well as improve the robustness of AI-driven detection methods through adversarial retraining and guardrails. When being used in critical scenarios with an extremely low error tolerance, combining AI-based attack detection with traditional detection methods is a promising direction to ensure reliability and robustness. For example, AI can categorize attack instances and generate insights from the data they’ve been trained on. Instead of manually analyzing every instance, security analysts can focus only on the representative cases from each category and extract static rules from AI.
@@
Facilitate triage and remediation. We first recommend building multi-agent systems for automated triage and patching. For example, we can integrate differential fuzzing to validate the correctness and security of AI-generated patches. We can also integrate formal verifications in patching agents to provide formal functionality and security guarantees. It is important to enable iterative refinements for these agents based on feedback. Such agents can also be leveraged to improve the automation and efficiency of SOC and security analysts’ workflow (e.g., data filtering and management). Second, we recommend training specific AI models to reason about complex program artifacts (e.g., atypical constructs), dependencies (inter-procedure dependencies and implicit constraints), and vulnerabilities. The methods discussed above about improving proactive testing and trained models can be applied here. We can also leverage the agent’s feedback as the reward signal for model training, such as training models to generate secure patches based on the PoC testing. Third, we can train specialized small models for planning or tool calling and use them to replace LLMs used in classification and patching agents. Fourth, we recommend involving more AI components in remediation deployments, such as testing case generation, agentic-based configuration, and dependency analysis. We also recommend building CI/CD pipelines that integrate these strategies to systemically reshape defense deployment. Finally, we recommend constructing more security-specific benchmarks for triage and remediation that encompass complex code bases across multiple programming languages.
@@
Improve the automation of defenses with provable guarantees. We believe frontier AI can significantly advance the scalability and automation of formal verification and call for increased efforts in this direction. First, we can use and fine-tune models to generate more complex proofs and constraints for theorem proving, as well as invariants for complex programs [314]. Second, we can use LLMs to generate necessary inputs for formal verification, such as program invariants and function summaries, where enabling verifiable generation with correctness guarantees is crucial. More broadly, secure code generation is a promising direction that can be used to turn unsafe latency code into safe counterparts. Early research has explored strategies like prefix learning [127] and constrained decoding [97]. We suggest focusing more on fundamental improvements in the models or agents, such as RL-based fine-tuning to improve models’ inherent capability of generating safe code. We can also design customized models and agents to transpile legacy programs into newer and memory-safe languages (e.g., C to Rust [87]). Third, AI can also improve constrained solvers. For example, SOTA solvers rely on somewhat random/rule-based mechanisms for backtracking and re-selecting verification paths to resolve conflicts. With frontier AI, solvers could learn more efficient strategies for making these critical decisions. Alongside building better AI-enhanced defenses, society must improve overall security posture and actively deploy these new defensive technologies.
7.3 Build Secure Hybrid Systems
@@
We call for more effort in developing secure mechanisms for hybrid systems, whether through secure-by-design and secure-by-construction approaches or by creating specialized proactive and reactive defenses. This is essential to address the significant gap between the rapid emergence of hybrid systems and the limited exploration of their security. First, the security community has developed various protection mechanisms for traditional symbolic systems, such as strong authentication, access control, and secure protocol designs. However, with the rise of hybrid systems, there is currently no standardized secure design framework or principles tailored for such systems. As such, we recommend designing secure system design frameworks and principles for hybrid systems. For example, these new frameworks and standards could adapt existing security enforcement mechanisms (e.g., privilege isolation), as well as create new mechanisms that are customized for the marginal risks unique to hybrid systems (e.g., tool permission controls and specific sandboxing). For example, having a more fine-grained rigorous isolation mechanism for AI systems also helps with safety and security, e.g., separating system prompts from user prompts helps mitigate prompt injections [51].
@@
Second, we recommend building real-time monitoring and alert mechanisms for hybrid systems. For example, we can extend existing AI attack detectors [9] to frontier AI, improve their efficiency, and continuously update them. In the long term, achieving provable security guarantees for hybrid systems is crucial. One possible direction is to follow a divide-and-conquer approach: decompose the verification of a hybrid system into sub-goals for its AI and traditional symbolic components, applying or extending specific techniques to address each. Overall, this remains an open challenge requiring greater attention and collaborative efforts from multiple computer science communities, including security, AI, programming languages, etc.
@@ 7.4 Recommendations for AI Development
Testing the cyberattack capabilities of new models before release and publishing detailed evaluation reports remains essential, as demonstrated by the GPT system card [140] and the Gemini dangerous capability evaluation [221]. As benchmarks evolve, AI developers should conduct more comprehensive testing beyond CTF competitions and collaborate with white-box hackers for real-world scenario testing. We can also explore implementing finer-grained privilege and access controls to ensure responsible use and avoid equivalence classes, enabling established defenders to access frontier AI more readily than attackers. However, implementation challenges remain, particularly in defining “established defenders.” While some arguments support open-source frontier AI to foster innovation, decisions about open-sourcing models and weights require careful risk assessment [153, 232].
Safety and security demand continued and increased attention. We recommend extending current comprehensive red-teaming practices more broadly across AI development, particularly for hybrid systems. For blue-teaming, we advocate a best-effort approach incorporating complementary strategies: model safety alignment, guardrails for models and agents, and system-level protections. Transparency remains a key limitation of frontier AI [232]. Considering the challenges in explaining complex model decisions, transparency in development processes is crucial for informed policymaking. Implementing varied transparency levels for different entities (public, trusted third parties, government) could balance information sharing risks and benefits, though this approach raises significant implementation challenges [232].
@@ 7.5 Mitigation of Human-Related Risks
To mitigate human-related attack risks, we first call for better user and developer education about AI attacks and security practices. Human awareness remains critical in cybersecurity, with research showing that developers’ security knowledge gaps and user unawareness significantly increase vulnerability and damage [20, 22]. Studies demonstrate AI can enhance educational efficiency across domains [50, 241, 226], suggesting the potential of leveraging frontier AI for security education. AI-driven personalized learning systems can deliver more efficient security training tailored to individual knowledge levels [264, 176]. Intelligent nudge mechanisms represent another effective approach [123, 163], where AI systems detect potential phishing or scams and provide timely alerts, enhancing real-time threat recognition while minimizing user fatigue.
However, education alone is insufficient, as users often remain the weakest link. As detailed in Sections 3.2 and 4.3, AI-powered attacks affect humans on a massive scale, while defensive measures lag behind rapidly evolving threats. We urgently need defensive research addressing frontier AI-enabled sophisticated attacks rather than focusing solely on traditional scenarios.
@@
LLM-enhanced malicious bot activities, including misinformation campaigns and fraud schemes, exemplify this challenge. AI-based protective measures are essential against these advanced threats [94]. Such mechanisms can analyze behavioral patterns, communication styles, and interaction frequencies to distinguish legitimate users from sophisticated bots. Platforms can implement intelligent screening mechanisms to filter out malicious bots attempting social engineering or fraud on social media.
Additionally, frontier AIs can create “honeypot” systems that attract and isolate these bots [230], engaging them in resource-draining interactions while learning their behaviors to develop better countermeasures. Given the rapid advancement of human-targeting attacks, game-theoretical approaches analyzing interactions between offensive and defensive frontier AI can provide valuable insights into this evolving security landscape.
8 Open Research Challenges and Questions

### v3

@@ 7.2 Enhance Cyber Defenses
We recommend enhancing traditional program analysis methods with AI-based rule extraction and planning, and developing specialized AI methods for security-specific challenges, like long-tail problems or attack shifts. It is also important to construct AI agents with flexible workflows and security-specific tools (e.g., static program graphs, dynamic testing), rather than focusing mainly on patching or fixed workflows. Within such frameworks, specialized reasoning models with stronger tool-calling abilities can be trained via SFT or RL. Finally, end-to-end defense benchmarks with executable environments are needed.
@@
Improve proactive testing. First, more efforts are needed in designing agent-based tools. Future efforts could focus on providing richer tool support, such as tools that help maintain access. With rich tools, we also need to design agents with more capabilities beyond the current ones, such as privilege escalation, compromising domain controllers, and lateral movements. Second, for static vulnerability detection, we recommend developing AI-based state pruning to alleviate state explosion and extracting static rules through model explanation to enhance knowledge bases. For fuzzing, one can explore efficient seed/mutator scheduling with RL and LLM-based grammar generation for domain-specific fuzzers. Third, we recommend fine-tuning customized models for better program state understanding and vulnerability reasoning, which requires training data with more program artifacts and context, especially dynamic information. Distillation of large reasoning models through supervised fine-tuning and reinforcement learning with proper reward functions is a promising direction. It is critical to scale AI-driven vulnerability detections at the repository level. Fine-tuning models with long context limits and building AI agents with program analysis tools (e.g., AST analysis and data/control flow analysis) could be promising directions. Finally, we recommend constructing vulnerability detection benchmarks that avoid data leakages in the testing set and have balanced vulnerability distributions that consider the long tails. Pairing human experts with AI can strike a balance between reducing manual efforts and improving data quality.
@@
Enhance attack detection. We first recommend constructing high-quality benchmarks that minimize the critical data and label issues, such as duplicated data, spurious correlations, biased samples, and wrong and missing labels. A recent research [15] provided a comprehensive summary of these issues. Given the challenges of collecting high-quality real-world attack data, collecting data from controlled environments [24] and creating synthetic data [247] could be efficient ways. Constructing hidden test sets with unseen attacks will help continuously retrain AI-driven defenses; once the currently hidden test sets are used for training, new test sets must be constructed and kept separate from the training data. Such practices have been applied in the natural language field: hiding test sets (e.g., SQuAD [186]), updating test sets with recent data (e.g., LiveBench [244]), or estimating and resolving training-test overlap [271].
@@
Although transformers address key challenges of other DNN models, it is important to develop customized learning methods for security-specific challenges, e.g., attack drifting and extremely imbalanced classes, as well as improve the robustness of AI-driven detection methods through adversarial retraining and guardrails. When being used in critical scenarios with an extremely low error tolerance, combining AI-based attack detection with traditional detection methods is a promising direction to ensure reliability and robustness. For example, AI can categorize attack instances and generate insights from the data they have been trained on. Instead of manually analyzing every instance, security analysts can focus only on the representative cases from each category and extract static rules from AI.
@@
Facilitate triage and remediation. We first recommend building multi-agent systems for automated triage and patching. For example, we can integrate differential fuzzing to validate the correctness and security of AI-generated patches. We can also integrate formal verifications in patching agents to provide formal functionality and security guarantees. It is important to enable iterative refinements for these agents based on feedback. Such agents can also be leveraged to improve the automation and efficiency of SOC and security analysts’ workflow (e.g., data filtering and management). Second, we recommend training specific AI models to reason about complex program artifacts (e.g., atypical constructs), dependencies (inter-procedure dependencies and implicit constraints), and vulnerabilities. The methods discussed above about improving proactive testing and trained models can be applied here. We can also leverage the agent’s feedback as the reward signal for model training, such as training models to generate secure patches based on the PoC testing. Third, we can train specialized small models for planning or tool calling and use them to replace LLMs used in agents, as general models may lack capabilities in calling domain-specific tools. This enables the models to learn unique agentic capabilities, as the data can only be obtained from the specific agents. Fourth, we recommend involving more AI components in remediation deployments, such as testing case generation, agent-based configuration, and dependency analysis. We also recommend building CI/CD pipelines that integrate these strategies to systemically reshape defense deployment. Finally, we recommend constructing more security-specific benchmarks for triage and remediation that encompass complex code bases across multiple programming languages.
@@
Improve the automation of defenses with provable guarantees. We believe frontier AI will significantly advance the scalability and automation of formal verification and call for increased efforts in this direction. First, we can use and fine-tune models to generate more complex proofs and constraints for theorem proving, as well as invariants for complex programs [258]. Second, we can use LLMs to generate necessary inputs for formal verification, such as program invariants and function summaries, where enabling verifiable generation with correctness guarantees is crucial. More broadly, secure code generation is a promising direction that can be used to turn unsafe latency code into safe counterparts. Early research has explored strategies like prefix learning [100] and constrained decoding [76]. We suggest focusing more on fundamental improvements in the models or agents, such as RL-based fine-tuning to improve their inherent capability of generating safe code. We can also design customized models and agents to transpile legacy programs into newer and memory-safe languages (e.g., C to Rust [67]). Third, AI can also improve constrained solvers. For example, SOTA solvers rely on random/rule-based mechanisms for backtracking and re-selecting verification paths to resolve conflicts. With frontier AI, solvers could learn more efficient strategies for making these critical decisions. Alongside building better AI-enhanced defenses, society must improve overall security posture and actively deploy these new defensive technologies.
Lesson and recommendations based on AIxCC. The AIxCC competition [1] highlights valuable insights into the current capabilities and limitations of AI-powered cybersecurity systems. The competition required teams to build systems to automatically analyze large codebases and generate PoCs and patches. Most systems in the final still heavily rely on traditional techniques for vulnerability detection and PoC generation. It shows that AI capabilities in these fundamental steps require significant improvement, which aligns with our study. In contrast, patch generation has shown significant evolution with heavy reliance on LLMs, which represent the SOTA in automated code generation. However, these systems still incorporate substantial human knowledge through carefully crafted workflows and prompts. Notably, team Theori [249] adopts a more agentic strategy by using a large number of agents across all steps, enabling them to achieve third place with simpler code structure and reduced human efforts.  These observations align with our recommendations on developing more advanced agents with agent-determined workflows and more program analysis tools, where agent-determined workflows can improve the generalizability and domain-specific tools can improve the agent’s capabilities in security tasks (e.g., static analysis-based context retrieval and fuzzing-based patch validation). In addition, the agentic systems can provide environment, data, and reward signals to continuously improve models’ capabilities in such systems.
7.3 Build Secure AI Agents
@@
We call for more effort in developing secure mechanisms for AI agents, whether through secure-by-design and secure-by-construction approaches or by creating specialized proactive and reactive defenses. This is essential to address the significant gap between the rapid emergence of AI agents and the limited exploration of their security. First, the security community has developed various protection mechanisms for traditional symbolic systems, such as strong authentication, access control, and secure protocol designs. However, with the rise of AI agents, there is currently no standardized secure design framework or principles tailored for such systems. This underscores the need for developing secure system design frameworks and principles for AI agents. These new frameworks and standards could adapt existing security enforcement mechanisms (e.g., privilege isolation), as well as create new mechanisms customized for the marginal risks unique to AI agents (e.g., tool permission controls and specific sandboxing). Having a more fine-grained rigorous isolation mechanism for AI systems also helps with safety and security, e.g., separating system prompts from user prompts helps mitigate prompt injections [38].
@@
Second, we recommend building real-time monitoring and alert mechanisms for AI agents. To this end, we can extend existing AI attack detectors [8] to frontier AI, improve their efficiency, and continuously update them. In the long term, achieving provable security guarantees for AI agents is crucial. One possible direction is to follow a divide-and-conquer approach: decompose the verification of an AI agent into sub-goals for its AI and traditional symbolic components, applying or extending specific techniques to address each. Overall, this remains an open challenge requiring greater attention and collaborative efforts from multiple communities, including security, AI, and programming languages.
@@ 7.4 Recommendations for AI Development
Testing the cyberattack capabilities of new models before release and publishing detailed evaluation reports remains essential [113, 180]. As benchmarks evolve, AI developers should conduct more comprehensive testing beyond CTF competitions and collaborate with white-box hackers for real-world scenario testing. We can also explore implementing finer-grained privilege and access controls to ensure responsible use and avoid equivalence classes, enabling established defenders to access frontier AI more readily than attackers. However, implementation challenges remain, particularly in defining “established defenders.” Decisions about open model access require careful risk assessment [122, 190].
Safety and security demand continued and increased attention. We recommend extending current comprehensive red-teaming and scaled eval practices more broadly across AI development, particularly for AI agents. For blue-teaming, we advocate a best-effort approach incorporating complementary strategies: model safety alignment, guardrails for models and agents, and system-level protections. Transparency remains a key limitation of frontier AI [190]. Considering the challenges in explaining complex model decisions, transparency throughout the entire AI lifecycle—including development, training, deployment, and use—is crucial for informed policymaking. Implementing varied transparency levels for different entities (public, trusted third parties, government) could balance information sharing risks and benefits, though this approach may introduce implementation challenges and potential conflicts among stakeholders [190].
@@ 7.5 Mitigation of Human-Related Risks
We first call for better user and developer education about AI attacks and security practices. Human awareness remains critical in cybersecurity, with research showing that developers’ security knowledge gaps and user unawareness significantly increase vulnerability and damage [17, 19]. Studies demonstrate AI can enhance educational efficiency across domains [37, 195, 185], suggesting the potential of leveraging frontier AI for security education. AI-driven personalized learning systems can deliver more efficient security training tailored to individual knowledge levels [215, 138]. Intelligent nudge mechanisms represent another effective approach [97, 129], where AI systems detect potential phishing or scams and provide timely alerts, enhancing real-time threat recognition while minimizing user fatigue. However, education alone is insufficient, as users often remain the weakest link. AI-powered attacks affect humans on a massive scale, while defensive measures lag behind rapidly evolving threats. We urgently need defensive research addressing frontier AI-enabled sophisticated attacks rather than focusing solely on traditional scenarios.
@@
LLM-enhanced malicious bot activities, including misinformation campaigns and fraud schemes, exemplify this challenge. AI-based protective measures are essential against these advanced threats [74]. Such mechanisms can analyze behavioral patterns, communication styles, and interaction frequencies to distinguish legitimate users from sophisticated bots. Platforms can implement intelligent screening mechanisms to filter out malicious bots attempting social engineering or fraud on social media. Frontier AIs can create “honeypot” systems that attract and isolate these bots [188], engaging them in resource-draining interactions while learning their behaviors to develop better countermeasures. Given the rapid advancement of human-targeting attacks, game-theoretically based attack and defense analysis can provide valuable insights into this evolving landscape.
8 Discussion and Conclusion
This paper presents an in-depth analysis of frontier AI’s current impacts on cybersecurity from three aspects (benchmarks, literature review, and expert survey), as well as projecting the future impact of AI and calls-to-action for steering AI to improve security. Even though current AI techniques have limitations, in this paper, we adopt an optimistic view: given the rapid evolution of models, AI holds significant potential to enhance the cybersecurity landscape. With this view, we provide recommendations on overcoming identified obstacles to promote AI in security. However, AI may fail, and there could be better alternative approaches. We leave the exploration of such alternatives for future work. This work highlights a few future directions. First, as a longitudinal study, we will continuously update our qualitative and quantitative analyses and expert survey to maintain an up-to-date understanding of AI’s evolving role in cybersecurity. Second, as AI is increasingly used in the physical world, future work can conduct a similar analysis for cyber-physical systems and broader hardware systems.
Ethical Considerations
First and foremost, this paper does not endorse or encourage the offensive use of AI. While our study examines the potential misuse of frontier AI systems and benchmarks their offensive capabilities, our intention is to inform the community and strengthen defenses. We acknowledge that research in this area carries inherent dual-use risks; insights that help defenders anticipate and mitigate attacks might also be misused by malicious actors. For this reason, we have exercised caution in the presentation of technical details and have focused on aggregated insights, high-level patterns, and defensive recommendations.
We believe that the benefits of publishing this work outweigh the risks. Specifically, we anticipate positive impacts by: (1) providing a systematic, big-picture view of the intersection of AI and cybersecurity; (2) identifying emerging risks before they are exploited at scale; and (3) offering actionable recommendations for researchers, practitioners, and policymakers to better prepare for the arrival of advanced AI systems in cybersecurity. Our hope is that this work will promote collaboration between the AI and security communities, leading to stronger standards, benchmarks, and defenses.
With respect to the expert survey conducted for this study, it was conducted between April and July 2025. All participants were clearly informed of its purpose, scope, and voluntary nature. They provided explicit consent for their responses to be used in aggregated analyses. No personally identifiable information was collected or reported. The study protocol was reviewed and approved by the Institutional Review Board (IRB) at UC Berkeley, ensuring that it meets established ethical standards for research involving human participants.
Finally, we emphasize that the broader community has a responsibility to handle this research responsibly. As AI capabilities advance rapidly, transparency, accountability, and continuous dialogue across technical, policy, and civil society domains will become critical in balancing benefits and risks.
Appendix A Performance of SOTA LLMs on Existing Attack and Defense Benchmarks
Table 5 and Table 6 summarize the performance of SOTA LLMs’ performance on popular attack and defense benchmarks discussed in the main text. In Table 5, all data are in percentage scale. The vulnerable code generation part reports the insecure code rates by instruction-based code generation. In the attack generation, we report the end-to-end attack success rate except for the RedCode-Gen which only involves code generation and we report the accuracy for it. For the SecCodePLT, we multiply the success rate of five steps to get the end-to-end success rate. For the CyBench and the NYU CTF benchmarks, we report the unguided solved rate and pass@1 rate separately. For the CyberBench benchmark, we report the average of the results in all sub-tasks. For the CyberMetric benchmark, we report the accuracy of the 2k questions. For the TACTL benchmark, we report the accuracy of the TACTL-183 benchmark. In Table 6, all data are in percentage scale. We report both autonomous (left) and assisted (right) success rates for the AutoPenBench. For the PrimeVul benchmark, we report the pair-wise correct prediction of the paired functions task with chain-of-thought. For the CRUXEval benchmark, we report the pass@1 rate on both input(left) and output(right) prediction results. For the SWE-bench-verified, we report the success rate. Note the agent architecture might be different.
Table 5: Performance of SOTA LLMs on existing attack benchmarks. All data are in percentage scale with the value range between 0 and 100. ’-’ means the model is not tested on the corresponding benchmark.
  Models Vulnerable code generation Attack generation CTF Cyber knowledge
CyberSecEval-3 SecCodePLT CyberSecEval-3 SecCodePLT RedCode-Gen RedCode-Exec CyBench NYU CyberBench CyberMetric TACTL
gpt-3.5 - - - - 32.5 - - - 62.6 88.10 -
gpt-4 35 - 40 - 66.9 64.50 - 7.00 69.9 91.00 -
gpt-4o - 44 - 0.2 72.5 77.23 12.5 9.50 - 91.25 85.2
GPT-4.5-preview - - - - - - 17.5 - - - -
o1-preview - - - - - - 10.0 - - - -
o3-mini - - - - - - 22.5 - - - -
claude-3.5-sonnet - - - 0 - 67.63 17.5 13.50 - - -
Claude 3.7 Sonnet - - - - - - 20 - - - -
gemini-1.5-pro - - - - - - 7.5 - - - -
gemini-1.0-pro - - - - - - - - - 84.00 -
deepseek-r1 - - - - - - - - - - 91.8
deepseek-v3 - - - - - - - - - - 86.3
llama-3.1-405b 39 - 49 - - - 7.5 - - - 88.5
llama-3.3-70b - - - - - - - - - - 78.7
llama-3.1-70b 35 47 41 0 - 76.70 5.0 - - - -
llama-3.1-8b - - - - - 62.87 - - - - -
llama-3-8b - - - - - - - - - 73.05 -
llama-2-7b - - - - 20.7 - - - 50.6 - -
llama-2-70b - - - - - - - - - 71.60 -
Table 6: Performance of SOTA LLMs on defense benchmarks. We report it only for the benchmarks for which results are publicly available. All data are in percentage scale with the value range between 0 and 100. ’-’ means the model is not tested on the benchmark. We report both autonomous (left) and assisted (right) success rates for the AutoPenBench.
  Models Pen. test Vul. detection PoC generation Patching
AutoPenBench PrimeVul CRUXEval SWE-bench-verified
gpt-3.5 - 6.21 49.1/63.3 0.4
gpt-4 - 12.94 74.8/81.9 22.4
gpt-4o 21.00/64.00 - - 38.8
o1 - - - 48.9
o3-mini - - - 49.3
claude-3.5-sonnet - - - 49.0
claude-3.7-sonnet - - - 70.3
deepseek-v3 - - - 42.0
deepseek-r1 - - - 49.2
code-llama-13b - - 39.1/39.3 -
code-llama-34b - - 50.4/46.0 -
Appendix B Expert Survey
B.1 Survey Questions
We present the survey questions and the number of responses we received for each item.
  • How effective do you think AI is in accelerating attacks against computer systems today? ($n=46$)
  • How effective do you think AI is in accelerating cybersecurity attacks against humans (e.g., social engineering attacks, deepfake, misinformation, etc) today? ($n=44$)
  • For attacks against systems, how much do you think each of the following attack activities would benefit from frontier AI in the near future (3-5 years)?
    • Scan the environment or harvest information (from social media) to identify potential targets ($n=21$)
    • Create attacks (malware) and encapsulate it as a weapon ($n=21$)
    • Transmit attacks (through e-mail, USB, web) ($n=26$)
    • Take actions to exploit certain vulnerabilities in the victim ($n=27$)
    • Install the malware on the victim’s system ($n=22$)
    • Establish a remote command channel, allowing attacker to access and manipulate the victim system ($n=27$)
    • Attacker achieves their ultimate objectives (e.g. data theft, sabotage) through the command and control channel ($n=21$)
  • For attacks against humans, how much do you think each attack method would benefit from frontier AI in the near future (3-5 years)?
    • Social engineering (e.g., Phishing email) ($n=23$)
    • Malicious content delivery (e.g., Malware download link) ($n=27$)
    • Impersonation (e.g., Profile cloning, deepfake) ($n=26$)
    • Platform and service fraud (e.g., False advertising, fake reviews) ($n=29$)
    • Information manipulation (e.g., Misinformation, deepfake) ($n=21$)
  • In what time frame do you believe the following attack activities against computer systems can be autonomously performed by frontier AI without human intervention besides providing necessary inputs? Consider technical feasibility and practicality for a general scenario that is relatively difficult, though it does not need to cover every possible setting.
    When answering the following questions, you can refer to the following setup: Imagine a network system with a web server and database, with internal users, and an active domain server. External users can interact with the network through the web server. All servers run Linux and may contain known N-day vulnerabilities such as Log4j and/or unknown 0-day vulnerabilities.
    • Craft phishing emails/websites/files, send them to users from an internal network, and trick them into clicking the malicious link ($n=22$)
    • Identify and locate the N-day and 0-day vulnerabilities in the vulnerable endpoint, e.g., web server ($n=25$)
    • Generate an attack based on the found vulnerabilities to gain access to the network through the web server, this will require designing and generating an exploit chain involving multiple vulnerabilities ($n=23$)
    • Establish the command and control channel after controlling the internal users and web servers ($n=17$)
    • Find a way to access the target data in the database, which requires gaining the necessary credentials from the active domain server ($n=23$)
    • Evade the possible defenses in the internal user’s system if blocked or detected ($n=21$)
    • Retrieve the target information and send it back to the attacker ($n=21$)
    • Conduct reverse engineering of complex software without any additional information other than stripped binaries (including disassembly and decompilation) ($n=20$)
  • In what time frame do you believe the following human-targeting attack activities can be autonomously performed by frontier AI without human intervention besides providing necessary inputs? Consider technical feasibility and practicality for a general scenario that is relatively difficult, though it does not need to cover every possible setting.
    • Execute sophisticated business email compromise by analyzing organizational communication patterns and mimicking writing styles, tone, and context to generate highly convincing impersonation emails ($n=24$)
    • Exploit deepfake technology to impersonate executives with the aim of manipulating employees into transferring funds or approving unauthorized contracts through falsified video or audio communications ($n=23$)
    • Analyze and replicate biometric patterns (gait, typing patterns, behavioral biometrics) to bypass multi-factor authentication systems protecting critical infrastructure ($n=23$)
  • How effective do you think frontier AI is in helping build defenses for computer systems (e.g., defense against MITER attacks, like malware and APT) today? ($n=45$)
  • How effective do you think frontier AI is in helping build defenses for humans (e.g., defense against human targeting attacks like deepfake and social bots) today? ($n=45$)
  • Regarding defenses for computer systems, how do you think each defense step would benefit more from frontier AI in the near future (3-5 years)?
    • Probe systems before an attack (e.g., pen-testing, fuzzing) to uncover vulnerabilities ($n=26$)
    • Identify malicious behavior as it happens and raise real-time alerts ($n=24$)
    • Reproduce the vulnerabilities and identify root causes ($n=26$)
    • Create patches or other counter-measures (isolation, blocking) ($n=29$)
    • Deploy and validate those fixes in production ($n=30$)
  • In what time frame do you believe the following defense activities for computer systems can be autonomously performed by frontier AI without human intervention besides providing necessary inputs? Consider technical feasibility and practicality for a general scenario that is relatively difficult, though it does not need to cover every possible setting.
    When answering the following questions, you can refer to the following setup: Imagine a network system with a web server and database, with internal users, and an active domain server. External users can interact with the network through the web server. All servers run Linux and may contain known N-day vulnerabilities such as Log4j and/or unknown 0-day vulnerabilities.
    • Detect and block phishing messages containing malicious payloads with a low false positive and false negative rate ($n=25$)
    • Detect most N-day vulnerabilities in the endpoints’ system and software with a low false positive and false negative rate ($n=18$)
    • Detect most 0-day vulnerabilities in the endpoints’ system and software with a low false positive and false negative rate ($n=21$)
    • Analyze the network pattern and build an intrusion detection system to monitor network traffic and detect malicious traffic from attackers (e.g., c2 patterns, extract sensitive data) with a low false positive and false negative rate ($n=23$)
    • Find the root causes for all vulnerabilities including N-day and 0-day ($n=21$)
    • Write patches that fix the vulnerabilities while maintaining existing functionalities ($n=19$)
    • Deploy patches, including generating configurations, analyzing dependencies, and resolving conflicts ($n=18$)
    • Develop and deploy necessary firewalls and sandboxing ($n=25$)
    • Write and deploy binary-level patches ($n=22$)
  • In what time frame do you believe the following defense activities for humans can be autonomously performed by frontier AI without human intervention besides providing necessary inputs? Consider technical feasibility and practicality for a general scenario that is relatively difficult, though it does not need to cover every possible setting.
    • Detect and prevent sophisticated business email compromise by analyzing communication patterns, authenticating sender behaviors, and identifying anomalies in writing styles, while automatically flagging suspicious emails and implementing real-time protection measures ($n=23$)
    • Analyze voice patterns, facial movements, and contextual inconsistencies in executive communications, enabling real-time detection of deepfake impersonation attempts ($n=22$)
    • Detect and block unauthorized access attempts that use synthetic or replicated biometric data by analyzing biometric patterns (gait, typing patterns, behavioral biometrics) in real-time ($n=27$)
  • How do you think frontier AI will impact the overall balance between attackers and defenders in cybersecurity? For each timeframe, assign a score from 0 to 100 points to indicate how much defenders will benefit from AI compared to attackers.
    100 points: Defenders exclusively gain the benefits from AI
    50 points: Both attackers and defenders equally benefit from AI
    0 points: Attackers exclusively gain the benefits from AI
    Consider both the technical feasibility and the likelihood of AI technologies being adopted by attackers and defenders in your assessment.
    • Within the next two years ($n=35$)
    • Within the next 5 years ($n=35$)
    • Within the next 10 years ($n=35$)
B.2 Demographic Information
In Table 7, we report the demographic information of the respondents who disclosed it.
Table 7: Participant demographics.
Category Demographic Count
Gender Female 7
Male 25
Non-binary 1
Age 18–24 6
25–34 15
35–44 5
45–54 4
55–64 2
Cybersecurity
Role
Industry 27
Academia 14
Non-profit 4
Government 4
Other 2
Cybersecurity
Experience
<< 1 year 9
1–3 years 13
4–6 years 4
7–10 years 3
>> 10 years 14
Familiarity with
LLMs/AI Agents
Expert 10
Very familiar 23
Moderately 10
Slightly 3
Not at all 1
B.3 Experts’ Views on the Future Balance Between Attackers and Defenders
Figure 4 shows how experts perceive the future balance between attackers and defenders. The results suggest that experts tend to believe the balance will tilt towards attackers in the near term, but will gradually narrow and reach parity over time.
(a) Within the next two years
(b) Within the next 5 years
(c) Within the next 10 years
Figure 4: Experts’ perceived balance between attackers and defenders in given timeframes. On the $x$-axis, a value of $100$ indicates that defenders exclusively gain the benefits from frontier AI, $50$ indicates parity, and 0 indicates that attackers exclusively gain the benefits from frontier AI.
Appendix C Open Research Challenges and Questions